OpenStudy (anonymous):

HELP with Quadratic Formula Word Problems!! One of the games at a carnival involves trying to ring a bell with a ball by h(t)= -16t + 39t. If the bell is 25 ft. above the ground, will it be hit by the bell?

OpenStudy (anonymous):

you said that this is a quadratic formula problem. Is one of the terms in the equation supposed to be a t^2 ("t squared") term? Right now, it looks like h(t) = -16t + 39t which is just h(t) = 23t... not quadratic.

OpenStudy (anonymous):

I was about to say that @JakeV8

OpenStudy (anonymous):

@JakeV8 Yeah it's most likely not quadratic. My teacher just titled the worksheet quadratic problems, I didn't know what to call this other than putting the problem

OpenStudy (anonymous):

Whoops sorry. wrong equation. it's h(t)= -16^2 + 39t.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@JakeV8 @breja

OpenStudy (anonymous):

Maybe like this? h(t) = -16t^2 + 39t

OpenStudy (anonymous):

oh okay wait one sec

OpenStudy (anonymous):

^^ Yeah it's that equation :) @JakeV8 sorry for the confusion!

OpenStudy (anonymous):

because the other wasn't quadratic lol

OpenStudy (anonymous):

h(t) = -16t2 + 39t dive both sides by t h = 39 + -16t simlify h = 39 + -16t

OpenStudy (anonymous):

@dellzasaur, I don't think that last post was quite correct... but sorry to disappear... my page froze up. h(t) = -16t^2 + 39t I'm pretty certain this means that the height of the ball is given by the function h(t). So, the question is, is there some value of "t" where the ball's height h(t) is equal to the bell's height of 25 feet.

OpenStudy (anonymous):

So, if the ball hits the bell, then the ball's height h(t) will be = 25. In other words, the question wants you to use the quadratic formula on the following: 25 = -16t^2 + 39t So you would start by putting the 25 on the same side as the rest... 0 = -16t^2 + 39t - 25 Now you can use quadratic formula: a = -16 b = 39 c = 25

OpenStudy (anonymous):

@JakeV8 The funtion is odd

OpenStudy (anonymous):

the function is odd

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

It is a function that is a parabola pointing down. This describes the motion of a ball thrown upward... imagine a softball thrown up high... it goes up, peaks at some height, then falls back to the ground.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I've done a problem like this before but..okay @JakeV8

OpenStudy (anonymous):

@breja, Earlier, when you said to divide both sides by "h", you just misunderstood the way they wrote the function. It wasn't "h times t" when they wrote "h(t)"... h(t) is a function, or a rule, that tells what the ball's height is given some time "t" since it was thrown.

OpenStudy (anonymous):

yea it was a mistake i read it wrong @JakeV8

OpenStudy (anonymous):

0 = -16t^2 + 39t - 25 quadratic formula: a = -16, b = 39, c = 25 Solutions to h(t) are: \[\frac{ -b \pm \sqrt{b^{2} -4ac} }{ 2a } = \frac{ -39 \pm \sqrt{39^{2} - 4(-16)(25)} }{ 2(-16) }\]

OpenStudy (anonymous):

simplify: \[\frac{ -39 \pm \sqrt{1521 - 1600} }{ (-32) } = \frac{-39 \pm \sqrt{-79}}{-32}\]

OpenStudy (anonymous):

ur so smart

OpenStudy (anonymous):

Uh oh :) That means the solutions "t" that make h(t) = 25 involve a square root of a negative number.... bad news!!! There are no real numbers that are the square root of a negative... That means there are no real number solutions that will allow h(t) to be equal to 25. And what that really means is that there are no real solutions where the ball's height is equal to the bell's height of 25 feet... In other words, no, the ball doesn't hit the bell. @dellzasaur -- finally finished this one :) @breja -- thank you :)

OpenStudy (anonymous):

@JakeV8 Thank you so much!!!! I got that too :)

OpenStudy (anonymous):

I tried @dellzasaur